OpenStudy (anonymous):

Convert the radian measure to degree measure. Use the value of π found on a calculator and round answers to two decimal places. 9pi/12

OpenStudy (bahrom7893):

pi is 180 degrees, so to convert, do this: \[\frac{9\pi}{12}*\frac{180}{\pi}=\frac{9*180}{12}=135\]

OpenStudy (anonymous):

thank you! i have 2 more questions i dont understand :(

OpenStudy (anonymous):

Convert the angle 13°38'35" to decimal degrees and round to the nearest hundredth of a degree.

OpenStudy (anonymous):

and this one Use the arc length formula and the given information to find s. Show your work for full credit. r = 15 ft θ = 35° s = ?

OpenStudy (bahrom7893):

how many minutes are in a degree?

OpenStudy (anonymous):

i dont know im lost this is new to me

OpenStudy (wolf1728):

60 minutes in a degree

OpenStudy (anonymous):

ok got it so like in an hour

OpenStudy (anonymous):

but i dont get how to get the answers

OpenStudy (jdoe0001):

to convert to Radians from angles given in Degrees => \(\bf\Large Radians = \cfrac{\textit{degrees} \times\pi}{180}\)

OpenStudy (anonymous):

so 13pi/180

OpenStudy (jdoe0001):

hhhmm,,ohh, you have radians, and you need degrees, ok

OpenStudy (jdoe0001):

9pi/12 I see

OpenStudy (anonymous):

wait what? :o

OpenStudy (jdoe0001):

well, isn't it what you have? \(\bf \cfrac{9\pi}{12}\)

OpenStudy (jdoe0001):

that's a Radian measure

OpenStudy (anonymous):

for which one?

OpenStudy (jdoe0001):

"Convert the radian measure to degree measure. Use the value of π found on a calculator and round answers to two decimal places. 9pi/12" ^ tis what you typed in

OpenStudy (jdoe0001):

to find the degrees unit for a Radian value then => \(\bf Degrees = \cfrac{\textit{radians}\times 180}{\pi}\)

OpenStudy (anonymous):

ok so what do i put in where?

OpenStudy (jdoe0001):

well, you were given \(\bf \cfrac{9\pi}{12}\) so you put that there :)

OpenStudy (anonymous):

wait which one was the degrees?? :o

OpenStudy (jdoe0001):

dunno, you'd need to get the value => \(\bf Degrees = \cfrac{\frac{9\pi}{12}\times 180}{\pi}\)

OpenStudy (anonymous):

Convert the angle 13°38'35" to decimal degrees and round to the nearest hundredth of a degree.

OpenStudy (anonymous):

and this one Use the arc length formula and the given information to find s. Show your work for full credit. r = 15 ft θ = 35° s = ?

terenzreignz (terenzreignz):

You need to work backwards with 13°38'35" Firstly, how many minutes are there, if you've got 35 seconds?

OpenStudy (anonymous):

38 minutes?

terenzreignz (terenzreignz):

No, I mean, 35 seconds, that's 7/12 of a minute, aye?

OpenStudy (anonymous):

oh ok sorry yea i got it

terenzreignz (terenzreignz):

So you've got 38 and 7/12 of a minute. Now divide \[\huge 38\frac7{12} = \frac{12(38)+7}{12}\] by 60, What do you get? Use your calculator...

OpenStudy (anonymous):

wait so you solve this equation and then how do you get 2 separate things to divide by 60?

terenzreignz (terenzreignz):

Okay, let's make things easier... to divide by 60, just multiply the denominator 12 by 60 \[\huge \frac{12(38)+7}{12\times60}\] Work this out...

OpenStudy (anonymous):

463/720

OpenStudy (anonymous):

??

terenzreignz (terenzreignz):

@jjdees lol yeah, but what is it in decimals?

OpenStudy (anonymous):

0.6431

terenzreignz (terenzreignz):

LOL just two decimal places as per your instructions ^_^

OpenStudy (anonymous):

oh right hehe my baad so 0.64

terenzreignz (terenzreignz):

Yup.. now attach that to the degree bit... 13 and that's your answer :P

OpenStudy (anonymous):

13.64?

terenzreignz (terenzreignz):

Voila

OpenStudy (anonymous):

I GOT IT THANNK YOU

OpenStudy (anonymous):

now how bout the other one hahah

terenzreignz (terenzreignz):

Arclength formula is \[\huge r \theta\] but \(\large \theta\) has to be in radians...

OpenStudy (anonymous):

ok this is so confusig where did that 0 with a line come from

terenzreignz (terenzreignz):

It's called 'theta' and you actually typed it yourself a while ago -.-

OpenStudy (anonymous):

well yea i copy pasted from google, and yes i googled a 0 with a line in it.

terenzreignz (terenzreignz):

Theta... It just means the angle-measure, lol Angles are usually represented by lowercase Greek letters, such as \[\large \alpha \qquad\beta\qquad \gamma \qquad \theta\]

OpenStudy (anonymous):

okey dokey

OpenStudy (anonymous):

so whats next

terenzreignz (terenzreignz):

That's it, just multiply the radius r with the angle measure \(\theta\) as long as it's in radians. What's your r-value?

OpenStudy (anonymous):

r=15ft 0thing=35 degrees

terenzreignz (terenzreignz):

Right, that 0thing (theta! you'll probably be more understood here that way) is equal to 35 degrees, now convert that to radians...

OpenStudy (anonymous):

hehe yeaaaa so i just multiply the two numbers?

terenzreignz (terenzreignz):

Nope... not that simple, you may only multiply the radius with the angle measure in radians. I already told you... convert 35 degrees to radians...

OpenStudy (anonymous):

argh, now how do you do that again?

terenzreignz (terenzreignz):

To convert form degrees to radians, multiply it by \[\huge \frac{\pi}{180^o}\]

OpenStudy (anonymous):

35pi/180

terenzreignz (terenzreignz):

lol yeah, that actually works :P And you multiply that to the radius, and that's your arclength XD

terenzreignz (terenzreignz):

you might want to simplify after that, but hey, that's the long and the short of it :P

OpenStudy (anonymous):

ok ok ok so to multiply the pie, do i actually get a decimal number or does it stay pi?

terenzreignz (terenzreignz):

I don't know, what does your question call for? In terms of pi? Or to some decimal place?

terenzreignz (terenzreignz):

You can always use the value \[\large \pi \approx 3.1415926535897932384626433832795028841...\]

terenzreignz (terenzreignz):

LOL yes, I needed that ^ ^_^

OpenStudy (anonymous):

doesnt say :(

terenzreignz (terenzreignz):

then just put it in terms of pi, but simplify , okay? I need to go now:P ----------------------------- Terence out